VIBRATIONAL SPECTRUM OF THE THIOMETHOXY (CH_{3}S) RADICAL INVESTIGATED WITH INFRARED-VACUUM ULTRAVIOLET PHOTOIONIZATION

\underline{HUI-LING HAN}, LUNG FU, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan.; YUAN-PERN LEE, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan..

We produced methylthio (or thiomethoxy, CH_{3}S) radicals by photodissociation of CH_{3}SH in a supersonic jet at 248 nm. The CH_{3}S^{+} ions were subsequently produced with the 1+1 IR-VUV photoionization and detected with the time-of-flight (TOF) technique. The IR spectrum of CH_{3}S was obtained on tuning the wavelength of the IR laser in the range 2780– 3280 cm^{−1} while monitoring the intensity of the CH_{3}S^{+} signal; the frequency of the VUV laser was maintained at 134.8 nm, 200 cm^{−1} below the ionization threshold of CH_{3}S (IE = 9.225 eV). This technique has an advantage over other IR-absorption techniques because its mass selectivity eliminates interferences from the precursor and other photolysis products such as H_{2}CS, CH_{3}, or CH_{3}SS. Absorption bands near 2820, 2904, and 3215 cm^{−1} were observed and tentatively assigned as transitions from the ground vibrational state to the 1^{1}, 4^{1} (a_{1}), and 5^{1}6^{2} states, respectively. These bands are in agreement with those reported for CH_{3}S produced via in situ photolysis of CH_{3}SH, CH_{3}SCH_{3}, and CH_{3}SSCH_{3} isolated in solid p-H_{2}.^{[1]} A new band near 2970 cm^{−1} that is consistent with that observed in photoelectron spectrum^{[2]} might be assigned to the transition from the ground vibrational state to the a_{1} component of the 5^{1}6^{2} state.

aM. Bahou and Y.-P. Lee, J. Chem. Phys. 133, 164316 (2010).

bR. L. Schwartz, G. E. Davico, and W. C. Lineberger, J. electron Spectrosc. Relat. Phenom. 108, 163 (2000).